Suppose you had as your personal energy source a quantity of matter and antimatter, which, when brought together, annihilated itself, that is, completely converted to radiant energy. What mass of fuel would be required in a lifetime of 81.8 years to provide power at a steady rate of 1.01 kW?
Answer #1

time t = 81.8 years = 2.579*10^9 s

power P = 1.01 * 10^3 W

Power = energy/time

P = mc^2/t

So the mass of fuel is

m = (Pt)/c^2

m = ( 1.01*10^3*2.579*10^9 )/(3.0*10^8)^2

m = 2.894*10^-5 kg